OpenStudy (anonymous):

What type of conic section is 4x^2-6x=-4y^2+y+1 A) circle B) ellipse C) can not be determined D) hyperbola

OpenStudy (savvy):

circle

OpenStudy (anonymous):

this is a different one. the last one was a circlr.....

OpenStudy (anonymous):

\[4x^2-6x=-4y^2+y+1 \]\[4x^2-6x+4y^2-y=1 \]\[x^2-3x/2+y^2-y/4=1/4\]Compete the squares, and you will get the standard form of a circle.

OpenStudy (anonymous):

so the conic section is a circle?

OpenStudy (anonymous):

In this case, yes. I'll continue where I left off last time....

OpenStudy (anonymous):

\[x^2−3x/2+y^2−y/4=1/4\]\[x^2−3x/2+9/16+y^2−y/4+1/64=1/4+9/16+1/64\]\[(x-3/4)^2+(y-1/8)^2=53/64\]

OpenStudy (anonymous):

Got it?